OpenStudy (anonymous):

Solve for x. Show your work. - 1/2 x < -12

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

multiply both sides by \(-2\) and flip the inequalty

OpenStudy (anonymous):

How would I go about flipping the inequality?

OpenStudy (nicolechoksy):

\[-1/2x < -12 \] Multiply both sides by -1 = \[1/2x > 12\] = \[x>12\div1/2\] = \[x>12\times2/1\] = \[x>24\]